Physicist Robert Adair, once appointed the "official physicist to the National League" by the commissioner of baseball, believes that the crack-of-the-bat can tell an outfielder how well the ball has been hit. According to Adair, a good hit makes a sound of 510Hz , while a poor hit produces a sound of 170Hz .

What is the difference in wavelength of these sounds? (in meters)
Answer #1

The basic equation is

sound speed = frequency x wavelength

So we have

wavelength = sound speed/frequency

The difference in two wavelengths will be

c/f1 - c/f2 = c(1/f1 - 1/f2)

Let sound speed = c = 343.4 m/s and we have

343.4(1/170 - 1/510)

= 343.4(0.0039215)

= 1.346666 meters